OpenStudy (anonymous):

5/6 - 4/12 Can someone show me step by step....PLEASE

OpenStudy (anonymous):

5/6 - 4/12 You must first reach a LCD (Least Common Denominator), which is 12 because 6 * 2 = 12. Multiply the first fraction by 2 on the top and bottom to create 10/12. 10/12 - 4/12 = 6/12. Simplify to 1/2.

OpenStudy (anonymous):

To subtract two fractions, first you need to find a common denominator

OpenStudy (anonymous):

I can..... find the LCD

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=5%2F6-4%2F12

OpenStudy (anonymous):

0.5

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf \frac{5}{6}-\frac{4}{12} = }\) \(\LARGE\color{blue}{ \bf \frac{5 \color{red} { \times2 } }{6\color{red} { \times2 } }-\frac{4}{12} = }\) \(\LARGE\color{blue}{ \bf \frac{10}{12}-\frac{4}{12} = }\) \(\LARGE\color{blue}{ \bf \frac{10-4}{12} = }\) \(\LARGE\color{blue}{ \bf \frac{6}{12} = }\) \(\LARGE\color{blue}{ \bf \frac{6 \color{red} { \div2 } }{12\color{red} { \div2 }} = }\) \(\LARGE\color{blue}{ \bf \frac{1}{2} ~.}\)

OpenStudy (solomonzelman):

just like in a math book :)

OpenStudy (anonymous):

Correction, just like wolframalpha.

OpenStudy (solomonzelman):

Yeah :) but wolframalpha only gives answers without explanations, unless you subscribe to wolframalpha PRO

OpenStudy (anonymous):

I need to make a living somehow, need someone to dust me every now and then.

OpenStudy (solomonzelman):

I use wolframalpha to check my work all the time, it is the best free solving program I've ever seen.

OpenStudy (anonymous):

Good job human!

OpenStudy (anonymous):

O_O wut about: https://www.mathway.com/

OpenStudy (solomonzelman):

you guys are to much into calculators. I don't have anything against calculators, but I think the best is to be able to solve in the head.

OpenStudy (anonymous):

Yeah. I agree. But when you have trig. and calculus... stuff gets real hard :/

OpenStudy (anonymous):

Please only use me to check your answers.

OpenStudy (solomonzelman):

lol

OpenStudy (anonymous):

hehehe XD